ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 112, 1999

An, J. The Alperin and Dade conjectures for the simple Conway’s third group

109

Buzzi, J. Markov extensions for multi-dimensional dynamical systems 357

Cohen, P. B. and A. Regev Asymptotics of multinomial sums and identities between multi-integrals

301

El Abdalaoui, E. H. La singularité mutuelle presque sûre du spectre des transformations d’Ornstein

135

Enriquez, B. and V. Rubtsov Quasi-Hopf algebras associated with 𝔰𝔩_{2} and complex curves

61

Hoffman, C. A loosely Bernoulli counterexample machine  237
Hoffman, C. A Markov random field which is K but not Bernoulli  249

Keen, L., J. R. Parker and C. Series Combinatorics of simple closed curves on the twice punctured torus

29

Lapid, E. M. Some results on multiplicities for SL(n)  157

Matoušek, J. and E. Matoušková A highly non-smooth norm on Hilbert space

1

Matoušková, E. See J. Matoušek

Mekler, A., A. Rosłanowski and S. Shelah On the p-rank of Ext 327

Montgomery, S. and H.-J. Schneider Prime ideals in Hopf Galois extensions

187

Newelski, L. Vaught’s conjecture for some meager groups  271

Parker, J. R. See L. Keen

Regev, A. See P. B. Cohen

Rosłanowski, A. See A. Mekler

Rubtsov, V. See B. Enriques

Schneider, H.-J. See S. Montgomery

Series, C. See L. Keen

Shelah, S. See A. Mekler